OpenStudy (anonymous):

Can someone please help me solve for x! 64^(-x)=4^(-3x)

zepdrix (zepdrix):

\[\Large\bf\sf \color{royalblue}{64\quad=\quad 4^3}\] Let's use this in our expression:\[\Large\bf\sf \left(\color{royalblue}{64}\right)^{-x}\quad=\quad 4^{-3x}\]

OpenStudy (jdoe0001):

\(\bf 64^{-x}=4^{-3x}\qquad {\color{blue}{ 64\implies 2^6\qquad 4\implies 2^2}} \\ \quad \\ 64^{-x}=4^{-3x}\implies (2^6)^{-x}=(2^2)^{-3x}\)

zepdrix (zepdrix):

Becomes:\[\Large\bf\sf \left(\color{royalblue}{4^3}\right)^{-x}\quad=\quad 4^{-3x}\]

OpenStudy (anonymous):

how do i figure out what x is equal to though?

zepdrix (zepdrix):

Using rules of exponents can be written as, \[\Large\bf\sf 4^{-3x}\quad=\quad 4^{-3x}\]The exponents multiply, remember your rule? :o

OpenStudy (anonymous):

no i can't remember it :(

OpenStudy (anonymous):

would x be 0?

OpenStudy (anonymous):

or would that mean that it is a true statement and i do not solve for x? :O

zepdrix (zepdrix):

Mmm Ya I guess 0 would work. 0 is a fairly trivial solution with exponential functions though :[

zepdrix (zepdrix):

Since the bases are now equal, it means the exponents have to be equal as well.\[\Large\bf\sf 4^{-3x}=4^{-3x}\qquad\implies\qquad -3x=-3x\]

zepdrix (zepdrix):

-3x=-3x, solve for x! :O

OpenStudy (anonymous):

well when i do that i get x=1, is that correct? :o

zepdrix (zepdrix):

Hmmm x=1 seems to also work for a solution. Hmmmmm what about x=2? Does that also hold true? :U

OpenStudy (anonymous):

yes i believe so :o so even though the question says to solve for x it wouldnt just be 1 or 0?

zepdrix (zepdrix):

Mmmm ok I'll stop being so sneaky :) lol Since the left and right side are completely identical, it means they will share EVERY VALUE OF X. If you wanted to `solve for x` though, \[\Large\bf\sf -3x=-3x\]We would add 3x to each side,\[\Large\bf\sf 0=0\]^This expression is true. Zero does in fact equal zero. So that tells us that the relationship is `true` for all x.

zepdrix (zepdrix):

If we had ended up with something like 0=1. That would tell us that the relationship is `false` for all x.

OpenStudy (anonymous):

oh i see! thank you very much! :) can you help me with another one similar to this? its just really hard :(

zepdrix (zepdrix):

Sure :U